[Brahms-l] Shift report 20040110 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 10 2004 - 00:07:29 EST
------------------------------------------------------------
Date: Sat Jan 10 00:07:28 2004, Shift: 16:00-24:00
Supervisor: KH, with: CEJ
Log book#15 pages 27 - 27

Shift Summary:
------------------------------------------------------------
1. The machine was perfoming well. Beam from 16.00 to 18.45
   and again from 22.10 (still running).

2. FS @ 4 deg. D1=3450A.
   MRS @ 90 deg. D5=1050B
   data runs: 
   9500, tr6 200K  
   9502, tr6  17K
   9505, tr6 120K
   9508, tr6 200K
   9509, tr6 200K
   9510 still running
   pedestal runs: 9506, 9507

3. TPC didn't have a good pedestal run before run 9500.
   H2 top 25 time is dead. T5-1 missing wires (since 9493).

      
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 10 00:07:40 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 10 2004 - 00:07:49 EST